OpenStudy (anonymous):

use the fact that the trigonometric functions are periodic to find the exact value of the given expression. do not use a calculator. tan 420

OpenStudy (anonymous):

tan 420 = tan (360 + 60) tan (x+y)= (tan x+tan y)/(1-tan x * tan y) tan(360+60) = (tan 360 + tan 60)/(1-tan 360 * tan 60) tan 360 = 0 tan 60 = sqrt(3) plug in values in above eqn =(0 + sqrt(3) ) / (1 - 0*sqrt(3)) =sqrt(3)/1 =sqrt(3)

OpenStudy (anonymous):

trigonometric functions are periodic so tan 420 = tan (420-360) = tan 60 = sqrt (3)